Sort by:[Date]

Map test procedure for the EPD

To test the mapping of the EPD, we change the vped setting in order to shift the pedestal value which can be picked up after reconstruction.  Additionally, we change the bias voltage so that we c

Rx crate pictures

 Attached are some pictures of the Rx crates.

Seven pattern files can resolve ALL ambiguity

In order to verify the tile mapping between "TUFF space" and "StEvent space", we have been setting Vped to different values  (in "TUFF space") and seeing the re

TMVA analysis

All TMVA related files can be found in dir

xeon-phi-dev:/star/u/kehw/TMVA/GenerateCandidateTree_SyncCuts

CalibRun15-Feb18

March 1 Update

New calibration notes

IV Scan Update

Executive Summary:  All dark currents look good prior to the start of the run.  Some useful code is attached.

Positronics connector - broken and fixed

Sorry for the brevity here. In February 2018, we saw a totally dead tile. We saw cosmics through the scope coming out of the Rx boards and eventually realized that the connector had a bent pin.

Zip-tie standoffs

 Two pictures of the zip-tie stand-offs used to create space between a fiber bundle and the East-side magnet busbar are attached.

Run12 pp500 Underlying Event Correction

run16 MTD bad module